OpenStudy (anonymous):

Solve for y. 2/3 y= 6

OpenStudy (anonymous):

y=9

OpenStudy (anonymous):

y = 9

OpenStudy (texaschic101):

2/3y = 6 --- divide both sides by 2/3 y = 6 / (2/3) y = 6 * 3/2 y = 18/2 y = 9 I at least wanted to show you how to do it.